OpenStudy (anonymous):

A test consists of multiple choice questions worth 2 points and essay questions worth 8 points. The test has 29 questions that are worth a total of 100 points. Use a model to determine how many multiple choice questions are on the test

OpenStudy (koikkara):

As per my Trial..... let x=no. of 2-point questions 29-x=no. of 8-point questions .. 2x+8(29-x)=100 2x+232-8x=100 6x=132 x=22 29-x=7 no. of 2-point questions=22 no. of 8-point questions=7 try whether u get the same.....